OpenStudy (anonymous):

|3x-4|+|9x-12|=12x x=??

OpenStudy (anonymous):

3x-4+9x-12=12x 0=-16 no solution

OpenStudy (anonymous):

oh so i can just collect like terms?

OpenStudy (anonymous):

when u collect them the x terms add up to zero

jimthompson5910 (jim_thompson5910):

The equation |3x-4|+|9x-12|=12x Breaks down into the following 4 equations (3x-4)+(9x-12)=12x or -(3x-4)+(9x-12)=12x or -(3x-4)-(9x-12)=12x or (3x-4)-(9x-12)=12x Solve each equation for x and remember to check your answers.

OpenStudy (anonymous):

do i collect like terms to solve each eqn for x?:S

jimthompson5910 (jim_thompson5910):

yes, you would solve each equation individually as you normally would with any linear equation.

OpenStudy (anonymous):

oh ok thanks